MKTEXLSR(1) MKTEXLSR(1) NAME mktexlsr−createls-Rdatabases SYNOPSIS mktexlsr[directories] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualKpathsea:Alibraryforpathsearching. mktexlsrisusedtogeneratethels-Rdatabasesusedbythekpathsealibrary.Itwillcreatethem forthespecifieddirectories,orforadefaultlistifnodirectoriesarespecified. OPTIONS mktexlsracceptsthefollowingoptions: --helpPrinthelpmessageandexit. --version Printversioninformationandexit. Kpathsea7.32x 4January1998 1